OpenStudy (anonymous):

Which of the following equations has no solution? 3x + 1 = 2x + 1 3x + 1 = 3x + 1 3x + 1 = 3x + 2

OpenStudy (anonymous):

i think its the second one....

OpenStudy (anonymous):

how????????????????????

OpenStudy (anonymous):

for no solution ..... \[a1/a2=b1/b2\neq c1/c2\]

OpenStudy (anonymous):

a1/a2= 3/3 = 1 b1/b2= 1/1= 1 c1/c2 = 0/0 =0 therefore a1/a2=b1/b2≠c1/c2

OpenStudy (anonymous):

undrstud?????? O.o

OpenStudy (anonymous):

no